# Speeding-up Large-scale LP Energy System Models: Using Graph-theory to Remove the Overhead Cost of Flexible Modeling

### v2

@@ Abstract
Energy system models are crucial for planning, supporting, and understanding energy transition pathways. Flexible energy modelling tools have emerged to provide practitioners, planners, and decision-makers with various alternatives to represent diverse energy systems, including green hydrogen or exclusively renewable-powered storage assets. The increased interaction between energy sectors, temporal resolution, and extensive geographical scopes have led to large-scale problems posing significant computational challenges. Despite improvements in computing power and linear programming (LP) solvers, large-scale LP models are often simplified, sacrificing fidelity to speed up solutions. This paper aims to debunk the misconception that an LP model’s simplicity cannot be improved without sacrificing fidelity. We propose exploiting the graph nature of energy systems using a single building block, the Energy Asset, to reduce computational complexity. By using only one building block, the Energy Asset, we avoid intermediary assets and connections, thus reducing the number of variables by 26% and constraints by 35%. This approach naturally speeds up solving times by 1.27 times without sacrificing model fidelity. Our illustrative case study demonstrates these improvements compared to traditional two-building-block approaches. This paper aims to raise awareness in the energy modelling community about the quality of LP models and shows that not all LPs are created equal. Our proposed method speeds up energy system models regardless of anticipated advances in software and hardware, allowing for the solution of larger and more detailed models with existing technology.
@@
[inst3]organization=Smart Energy and Built Environment, VTT Technical Research Centre of Finland Ltd.,country=Finland
@@
Fewer variables and constraints maintain accuracy in energy optimisation models.
Network graph concept reduces variables and speeds up model creation and solving.
Benefits increase with problem size, aiding large-scale energy system modelling.
Flexible models allow diverse approaches impacting model performance.
Conscious decision-making is crucial in representing energy systems effectively
@@ 1.1 Motivation
Although currently available models can integrate the energy sector coupling in the models in different ways, the implications of having flexibility in the modelling choices have not been given much attention. Computational performance is affected by these modelling choices, especially in large-scale energy problems like those in regional-level studies (e.g., European case studies). As the energy modelling community is increasingly exploring new system configurations to integrate different processes, such as green hydrogen production, hybrid operation between a storage asset and a renewable asset, or small modular reactors producing both electricity and heat, there is also a need for general and flexible modelling structures that can be adapted to new configurations without compromising the model’s performance. Therefore, it is important to identify the computational implications of different modelling choices in this new context.
@@ 1.2 Solving methods in energy systems modelling
Many researchers view Linear Programming (LP) as the simplest problem representation that cannot be improved without sacrificing accuracy in energy systems modelling. This is a common misconception, which can lead to the creation of large overhead model sizes. There are at least two alternatives to improve the representation without losing accuracy. First, LP models can be reformulated to leverage specific characteristics, such as sparsity, or tailored for specialised algorithms, thus enhancing computational efficiency without sacrificing accuracy. For example, specialised algorithms like the Hungarian Algorithm for the assignment problem [1] and the Push-Relabel method for network flow problems [2] often perform better than general LP solvers. These cases suggest that rather than relying solely on advancements in LP solvers or hardware, researchers should explore specialised methods to improve efficiency. Second, since researchers widely use LP models, we explore another alternative to improve the formulation that still uses LP: reformulating the problem to exploit the graph nature of energy system models. All models are not the same, and improving the quality of an LP model means that it retains fidelity while solving faster. Hence, proposing a reformulation that can result in more significant advantages as the problem size increases is especially important for large-scale energy problems.
1.3 Building blocks for flexible modelling
@@
Multi-physics energy system models can be formulated in many ways, ranging from process-specific equations to a more generic approach where concepts such as nodes and units represent a wide variety of conversion, production, consumption, transfer, and storage processes [3]. In the process-specific approach, each process is described by its own equation or set of equations. Models may have been built this way due to historical reasons: at first the intention was to model a specific sector, such as the power system, and only afterwards was the model expanded to other energy sectors. An example is the early open-source model Balmorel [4], which was initially a power system model but has since been expanded to most energy sectors. This trend has also been seen in commercial options like Plexos [5]. Another example is the COMPETES-TNO model [6] that has incorporated hydrogen sector-specific constraints to a power system model. The trend has been driven by the need to model decarbonization pathways involving several sector coupling technologies, e.g., hydrogen production by electrolysis, electric vehicles, and heat production from electricity [7]. The main drawback of the process-specific approach is that the number of possible interactions between energy system components can become very large, making the model unwieldy to maintain and expand. Besides, there are only a limited number of different ways processes can be described in linear programmes (LPs) or mixed-integer linear programmes (MIPs).
@@
Another strategy is to formulate methods between higher-level concepts like nodes, connections and units that act as building blocks (BB) to build a more general/generic energy system. Each BB can then choose an appropriate method for every particular process. This limits the number of formulations to the number of supported methods; however, it provides flexibility to the user on the modelling options. It also means that users can add new process types without writing new code, provided that an appropriate method is available for them in the model. This approach has been used by models like the IRENA FlexTool [8] and, to an extent, SpineOpt [9].
@@
A further perspective is that energy system models can be seen as network graphs that illustrate the connections among various energy assets in different sectors or energy carriers [10]. From this viewpoint, different types of nodes typically serve as the basic building blocks that link all the elements such as producers, converters, storage, and consumers. However, existing literature has not delved into the potential use of the natural graph structure of energy systems to directly link the assets in order to improve the computational efficiency of models.
@@ 1.4 Contribution
We propose to exploit the graph nature of energy systems and replace the traditional BB, nodes, units and connections with only one BB: Energy Assets (vertices) and use energy flows (arcs) to connect them to each other. Thus, it inherently avoids unnecessary extra constraints and variables required by different building blocks while keeping the full model flexible and speeding up solving times without sacrificing the model’s fidelity.
@@ 2 Quality of LP models
A common misconception is that an LP is the simplest representation of a problem, which cannot be improved without sacrificing its fidelity. That is, the only way to speed up an LP without losing accuracy is through improvements in computing power (hardware) and LP solvers (software). This belief can lead modellers to inadvertently create large overhead model sizes, assuming the model is very efficient since LP is the most simplified you can go. When the model size is still potentially problematic, the common belief is that the only other option to speed up large-scale LPs is to solve an (over)simplified, smaller model, which sacrifices its fidelity.
@@
Although discussing model quality in LP Models is not common, model quality is a well-known concept in mixed-integer programming (MIP). The quality of an MIP model is defined by its tightness, that is, how near is its relaxed LP feasible region to that of the integer one. The tightest possible model (convex hull) can solve an MIP as an LP, greatly lowering the computational burden. However, trying to tighten an MIP formulation often implies increasing its size hence there is a trade-off between the tightness and compactness of an MIP.
@@
What then defines the model quality of an LP model? The quality of an LP model is defined by its size. That is, a more compact model, i.e., less constraints/variables and nonzeros, has higher quality than a less compact one with the same fidelity, and hence it is expected to solve faster. Here, we differentiate the model quality with numerical-related issues, which can also slow down solving times. That is, model quality is independent from the data used. Of course, the modeller should be careful when populating the model with data to avoid LP numerical issues, such as degeneracy, numerical stability and ill-conditioning [11].
@@
How to improve the quality of an LP model? how to lower its size without sacrificing its fidelity? To the best of our knowledge, there are three ways to lower one of the dimensions of a problem. First, the trivial option is to remove equalities. Each equality means that a variable in that equality can be removed by replacing its equivalent in the other constraints. Although this procedure lower the number of variables, it increases the number of nonzeros. Second, the Fourier–Motzkin procedure eliminates a set of variables, creating another model such that both models have the same solutions over the remaining variables. Fourier–Motzkin procedure could also eliminate constraints if applied on the dual of the formulation. However, this procedure comes at the expense of producing, an often, exponential number of constraints and non-zeros, thus creating worse bottlenecks, slowing down LP solving times. The third way of LP reformulation involves splitting dense columns into sparser ones [12]. Although this procedure increases the number of constraints and variables, it could speeup solving times when the number of nonzeros is the bottle-neck.
@@
Lowering the size of an LP model is not a trivial task, since lowering one dimension comes at the expense of sacrificing another dimension, which can become a new bottleneck potentially damaging the quality of the model instead of improving it. In this paper, we propose a change of paradigm for flexible energy system modelling, we propose using a single building block, the energy asset, thus fully exploiting the network nature of the system. This proposed reformulation lowers all three dimensions of the problem simultaneously, the number of variables, constraints and non-zeros. This higher quality LP reformulation naturally lowers both model creation times and solving times, obtaining more significant advantages as the problem size increases.
3 Building blocks for flexible modelling
@@
Energy system modelling requires the use of building blocks to describe the connectivity of different elements. One such block is the energy asset, which includes elements that can produce, consume, convert, or store energy. Current state-of-the-art energy systems models use at least two or three building blocks, such as nodes, connections and units. Energy system models often use nodes as an additional building block to connect energy assets and establish energy balances [13]. Some models have a storage option in their node balance [8] while others maintain a separate building block for it [14]. Moreover, some models include an extra building block to describe the connections between nodes [9]. Although having multiple building blocks to model an energy system seems appealing due to the flexibility to represent various configurations, it often comes with a computational cost that becomes more significant for large-scale problems.
@@
In this paper, we propose a different approach using only one building block: energy assets, represented as the vertices of a graph, with the flows between the assets as the arcs. Adopting this approach makes the node concept irrelevant since we have a direct connection between assets. This approach simplifies the graph structure of the energy system, allowing for flexible modelling options that reduce the number of variables and constraints required to model the same problem. This reduction has computational benefits in both creating the optimisation model and the time to solve it, as demonstrated in the experiments discussed in Section 5.
@@
Having flexible modelling options is becoming more relevant in energy system analysis with the appearance of new configurations, such as the hybrid configurations of storage and renewable assets. Figure 1 shows an example of such hybrid configuration, where storage can only be charged from solar PV but the storage and solar PV assets can both deliver energy to the grid. We can take this use case as an example to understand the difference between various approaches. In this case, the storage asset can only be charged from the renewable asset and not from the grid. Figure 1 gives an overview of each modelling approach for this example. In the following sections, we describe each approach in detail. It is worth noting that different modelling approaches represent the same situation, i.e., they are all modelling exactly the same problem, and this example is not unique. For instance, green hydrogen production is another example where a conversion asset, such as an electrolyser, can only produce hydrogen from the renewable asset; however, the renewable asset can still send energy to the network.
3.1 Three building blocks with four flow variables (3BB-4F)
@@
This approach uses three BB: energy assets, balance nodes, and connections to link the nodes. In addition, the connection uses four positive variables to link the nodes, two for each node it connects to. This approach is advantageous as it can represent situations where the incoming flow might differ from the outgoing flow, which are needed to model transmission or distribution networks losses or gas flows with linepack. However, when modelling network constraints without losses or differences between the incoming and outgoing flow, it creates extra variables and constraints.
@@
Figure 1(a) shows the battery (bt) with a renewable (pv) example for the 2BB-4F. The connection point (cp) is an auxiliary node connecting the two assets. The electricity demand (ed) is balanced in an extra node. Moreover, the maximum capacity of the flow coming from the connection line (cl) to the connection point must be zero to avoid charging the battery from the grid (i.e., electricity demand balance node). Therefore, we need 8 variables (7 flow variables + 1 storage level ) and 11 constraints (8 capacity limits + 2 node balances + 1 storage balance).
@@
: flow from the battery to the connection point (i.e., battery discharge) with a maximum capacity
: flow from the connection point to the battery (i.e., battery charge) with a maximum capacity
: flow from the solar pv to the connection point with a maximum capacity
@@
3.2 Two building blocks with two flow variables (2BB-2F)
This approach uses as building blocks the energy assets and the balance nodes. In addition, it uses two positive flow variables to connect the nodes instead of the four variables used in the previous approach. The two flow variables are needed to model transmission or distribution networks losses. However, gas flows with linepack cannot be modelled correctly with only two flow variables. Depending on the input data, the incoming and outgoing flow can have values simultaneously, especially when losses are not considered. In such cases, modellers typically use the net value as the transfer between the nodes. Alternatively, they may include a binary variable to prevent simultaneous incoming and outgoing flows, although this would convert the problem into a Mixed-Integer Programming problem.

### v3

# Debunking the Speed-Fidelity Trade-Off: Speeding-up Large-Scale Energy Models while Keeping Fidelity

@@ Abstract
Energy system models are essential for planning and supporting the energy transition. However, increasing temporal, spatial, and sectoral resolutions have led to large-scale linear programming (LP) models that are often (over)simplified to remain computationally tractable—frequently at the expense of model fidelity. This paper challenges the common belief that LP formulations cannot be improved without sacrificing their accuracy. Inspired by graph theory, we propose to model energy systems using energy assets (vertices), as a single building-block, and flows to connect between them. This reduces the need for additional components such as nodes and connections. The resulting formulation is more compact, without sacrificing accuracy, and leverages the inherent graph structure of energy systems. To evaluate performance, we implemented and compared four common modelling approaches varying in their use of building blocks and flow representations. We conducted experiments using TulipaEnergyModel.jl and applied them to a multi-sector case study with varying problem sizes. Results show that our single-building-block (1BB-1F) approach reduces variables and constraints by 26% and 35%, respectively, and achieves a 1.27x average speedup in solving time without any loss in model fidelity. The speedup increases with problem size, making this approach particularly advantageous for large-scale models. Our findings demonstrate that not all LPs are equal in quality and that better reformulations can lead to substantial computational benefits. This paper also aims to raise awareness of model quality considerations in energy system optimisation and promote more efficient formulations without compromising fidelity.
@@
[inst3]organization=Design and Operation of Energy Systems, VTT Technical Research Centre of Finland Ltd.,country=Finland
@@
[inst5]organization=Nodal-Tools Ltd.,country=Finland
The proposed single modelling block reduces size, speeding up creation and runtime
Fewer variables and constraints without losing accuracy in energy optimisation models
Benefits increase with problem size, aiding large-scale energy system modelling
Enhancing modelling with a more direct and flexible connection between assets
@@ 1.1 Motivation
Although currently available models can integrate the energy sector coupling in the models in different ways, the implications of having flexibility in the modelling choices have not been given much attention. Computational performance is affected by these modelling choices, especially in large-scale energy problems like those in regional-level studies (e.g., European case studies) for long-term expansion planning (e.g., pathway planning). As the energy modelling community is increasingly exploring new system configurations to integrate different processes, such as green hydrogen production, hybrid operation between a storage asset and a renewable asset, or small modular reactors producing both electricity and heat, there is also a need for general and flexible modelling structures that can be adapted to new configurations without compromising the model’s performance. Therefore, it is important to identify the computational implications of different modelling choices in this new context.
@@ 1.2 Solving methods in energy systems modelling
Many researchers view Linear Programming (LP) as the simplest problem representation that cannot be improved without sacrificing accuracy in energy systems modelling. This is a common misconception, which can lead to the creation of large overhead model sizes. There are at least two alternatives to improve the representation without losing accuracy. First, LP models can be reformulated to leverage specific characteristics, such as sparsity, or tailored for specialised algorithms, thus enhancing computational efficiency without sacrificing accuracy. For example, specialised algorithms like the Hungarian Algorithm for the assignment problem [1] and the Push-Relabel method for network flow problems [2] often perform better than general LP solvers. These cases suggest that researchers should explore specialised methods to improve efficiency rather than relying solely on advancements in LP solvers or hardware. Second, since researchers widely use LP models, we explore another alternative to improve the formulation that still uses LP: reformulating the problem to exploit the graph nature of energy system models. All models are not the same, and improving the quality of an LP model means that it retains fidelity while solving faster. Hence, proposing a reformulation that can result in more significant advantages as the problem size increases is especially important for large-scale energy problems.
1.3 Flexible modelling in energy system models
@@
Multi-physics energy system models can be formulated in many ways, ranging from process-specific equations to a more generic approach where concepts such as nodes and units represent a wide variety of conversion, production, consumption, transfer, and storage processes [3]. In the process-specific approach, each process is described by its equation or set of equations. Models may have been built this way for historical reasons: at first the intention was to model a specific sector, such as the power system, and only afterwards was the model expanded to other energy sectors. An example is the early open-source model Balmorel [4], which was initially a power system model but has since been expanded to most energy sectors. This trend has also been seen in commercial options like Plexos [5]. Another example is the COMPETES-TNO model [6] that has incorporated hydrogen sector-specific constraints to a power system model. The trend has been driven by the need to model decarbonization pathways involving several sector coupling technologies, e.g., hydrogen production by electrolysis, electric vehicles, and heat production from electricity [7]. The main drawback of the process-specific approach is that the number of possible interactions between energy system components can become very large, making the model unwieldy to maintain and expand. Besides, there are only a limited number of ways processes can be described in linear programmes (LPs) or mixed-integer linear programmes (MIPs).
@@
Another strategy is to formulate methods between higher-level concepts like nodes, connections and units that act as building blocks (BB) to build a more general/generic energy system. Each BB can then choose an appropriate method for every particular process. This limits the number of formulations to the number of supported methods; however, it provides flexibility to the user on the modelling options. It also means that users can add new process types without writing new code, provided an appropriate method is available in the model. This approach has been used by models like the IRENA FlexTool [8] and, to an extent, SpineOpt [9].
@@
A further perspective is that energy system models can be seen as network graphs that illustrate the connections among various energy assets in different sectors or energy carriers [10]. From this viewpoint, different types of nodes typically serve as the basic building blocks that link all the elements, such as producers, converters, storage, and consumers. However, the existing literature has not delved into the potential use of the natural graph structure of energy systems to link the assets directly in order to improve the computational efficiency of models.
@@ 1.4 Contribution
We propose the Energy Asset as a single building block, allowing a more direct connection between components by leveraging the graph structure of energy systems. The proposed strategy replaces the traditional building block, nodes, units and connections with only one BB: Energy Assets (vertices) and use energy flows (arcs) to connect them. Thus, it inherently avoids unnecessary elements in between with their associated extra constraints and variables, while keeping the full model flexible and speeding up solving times without sacrificing its fidelity.
@@ 2 Quality of LP models
A common misconception is that an LP is the simplest representation of a problem for a desired fidelity, and that the model cannot be simplified without compromising the value of its solution [11]. That is, the only way to speed up an LP without losing accuracy is through improvements in computing power (hardware) and LP solvers (software). This belief can lead modellers to inadvertently create large overhead model sizes, assuming the model is very efficient since LP is the most simplified you can go. When the model size is still potentially problematic, the common belief is that the only other option to speed up large-scale LPs is to solve an (over)simplified, smaller model, which sacrifices its fidelity.
@@
Although discussing model quality in LP Models is not common, model quality is a well-known concept in mixed-integer programming (MIP). The quality of an MIP model is defined by its tightness, that is, how near is its relaxed LP feasible region to that of the integer one. The tightest possible model (convex hull) can solve an MIP as an LP, greatly lowering the computational burden. However, trying to tighten an MIP formulation often implies increasing its size; hence, there is a trade-off between the tightness and compactness of an MIP.
@@
What then defines the model quality of an LP model? The quality of an LP model is defined by its size. That is, a more compact model, i.e., fewer constraints/variables and non-zeros, has higher quality than a less compact one with the same fidelity, and hence, it is expected to solve faster. Here, we differentiate the model quality with numerical-related issues, which can also slow down solving times. That is, model quality is independent of the data used. Of course, the modeller should be careful when populating the model with data to avoid LP numerical issues, such as degeneracy, numerical stability and ill-conditioning [11].
@@
How can the quality of an LP model be improved? How can it be lowered in size without sacrificing its fidelity? To the best of our knowledge, there are three ways to lower one of the dimensions of a problem. First, the trivial option is to remove equalities. Each equality means that a variable in that equality can be removed by replacing its equivalent in the other constraints. Although this procedure lowers the number of variables, it increases the number of non-zeros. Second, the Fourier–Motzkin procedure eliminates a set of variables, creating another model in which both models have the same solutions over the remaining variables. Fourier–Motzkin procedure could also eliminate constraints if applied to the dual formulation. However, this procedure comes at the expense of producing an often exponential number of constraints and non-zeros, thus creating worse bottlenecks and slowing down LP-solving times. The third way of LP reformulation involves splitting dense columns into sparser ones [12]. Although this procedure increases the number of constraints and variables, it could speed up solving times when the number of non-zeros is the bottleneck.
@@
Lowering the size of an LP model is not a trivial task since lowering one dimension comes at the expense of sacrificing another dimension, which can become a new bottleneck, potentially damaging the quality of the model instead of improving it. In this paper, we emphasise in a change of paradigm for flexible energy system modelling; we use a single building block, the energy asset, thus fully exploiting the network nature of the system. This reformulation simultaneously lowers all three dimensions of the problem: including the number of variables, constraints,, and non-zeros. This higher-quality LP reformulation naturally lowers both model creation and solving times, obtaining more significant advantages as the problem size increases.
3 Energy systems modelling and building blocks for flexible modelling
Energy systems can be represented as directed graphs, where the vertices often denote balancing nodes, and the edges represent the energy flows between these components. In this paper, we denote the graph’s vertices as the energy system’s Building Blocks (BB). Depending on their characteristics in the energy system, these building blocks represent Energy Assets that can produce, consume, convert, store, or balance energy.
@@ 3 Energy systems modelling and building blocks for flexible modelling
Current state-of-the-art energy systems models use at least two or three building blocks, such as nodes, connections and units. Energy system models often use nodes as an additional building block to connect energy assets and establish energy balances [13]. Some models have a storage option in their node balance [8], while others maintain a separate building block for it [14]. Moreover, some models include an extra building block to describe the connections between nodes [9]. Although having multiple building blocks to model an energy system seems appealing due to the flexibility to represent various configurations, it often comes with a computational cost that becomes more significant for large-scale problems.
@@
In this paper, we highlight the advantages of using only one building block: energy assets, represented as the vertices of a graph, with the flows between the assets as the arcs. Adopting this approach makes the node concept irrelevant since we directly connect assets. This approach simplifies the graph structure of the energy system, allowing for flexible modelling options that reduce the number of variables and constraints required to model the same problem. This reduction has computational benefits in both creating the optimisation model and the time to solve it, as demonstrated in the experiments discussed in Section 5.
3.1 Generic formulation
While each approach may have more specific constraints and differing notations, we can generalize them all by using the proposed energy asset as the only building block. The following formulation makes comparing the formulations in the following sections more straightforward.
Where:
: set of all assets with elements
: subset of investable assets
: subset of consumer assets
: subset of storage assets
: subset of conversion assets
: set of flows between two assets with elements
: set of all timesteps with elements
: investment unit variable of asset
: flow variable from asset to asset in timestep
: storage level variable of asset in timestep
: investment cost parameter of asset
: operation cost parameter of flow between two assets
: demand parameter of the consumer asset in timestep
: initial storage level parameter of the storage asset
: initial number of units parameter of the asset
: maximum capacity parameter of the asset
, : maximum flow between two assets in both directions.
: maximum storage capacity parameter of the storage asset
: maximum investment limit parameter of the asset
: efficiency parameter of flows going into the asset
: efficiency parameter of flows going out the asset
The objective function (1) minimises the investment cost and the operation variable cost of the flow between two assets. Equations (2), (3) and (4) define the balance constraint for consumer, storage, and conversion assets. The capacity limit of the flows (5) is defined by the maximum capacity of the asset, while the maximum charging limit for storage assets is defined by (6). Equation (7) represents the maximum storage capacity, and (8) defines the bounds for the flow variable. Equation (9) and (10) limit the availability of the investment variables for the assets. Annex A shows how to model a DC power flow and unit commitment in the proposed approach. In addition, the TulipaEnergyModel [15] documentation111https://tulipaenergy.github.io/TulipaEnergyModel.jl/dev/40-formulation/ shows a comprehensive overview of the proposed approach.
3.2 Illustrative example
@@
Having flexible modelling options is becoming more relevant in energy system analysis with the appearance of new configurations, such as the hybrid configurations of storage and renewable assets. Figure 1 shows an example of such a hybrid configuration, where storage can only be charged from solar PV. Still, the storage and solar PV assets can both deliver energy to the grid. We can take this use case as an example to understand the difference between various approaches, which only considers operation variables for the sake of simplicity . In this case, the storage asset can only be charged from the renewable asset and not from the grid. Figure 1 gives an overview of each modelling approach for this example. In the following sections, we describe each approach in detail. It is worth noting that different modelling approaches represent the same situation, i.e., they are all modelling exactly the same problem, and this example is not unique. For instance, green hydrogen production is another example where a conversion asset, such as an electrolyser, can only produce hydrogen from the renewable asset; however, the renewable asset can still send energy to the network.
3.2.1 Three building blocks with four flow variables (3BB-4F)
@@
This approach uses three BB: energy assets, balance nodes, and connections to link the nodes. In addition, the connection uses four positive variables to link the nodes, two for each node it connects to. This approach is advantageous as it can represent situations where the incoming flow might differ from the outgoing flow, which is needed to model transmission or distribution network losses or gas flows with linepack. However, modelling network constraints without losses or differences between the incoming and outgoing flow creates extra variables and constraints.
@@
Figure 1(a) shows the battery (bt) with a renewable (pv) example for the 2BB-4F. The connection point (cp) is an auxiliary node connecting the two assets. The electricity demand (ed) is balanced in an extra node. Moreover, the maximum capacity of the flow coming from the connection line (cl) to the connection point must be zero to avoid charging the battery from the grid (i.e., electricity demand balance node). Therefore, we need 8 variables (7 flow variables + 1 storage level ) and 11 constraints (8 capacity limits + 2 node balances + 1 storage balance). The index is not included in the following equations for simplicity.
@@
: flow from the battery to the connection point (i.e., battery discharge) limited by battery capacity
: flow from the connection point to the battery (i.e., battery charge) limited by battery capacity 222Charging and discharging capacities may vary based on the type of storage asset; for simplicity, we consider them equal.
: flow from the solar pv to the connection point limited by solar PV availability
@@
3.2.2 Two building blocks with two flow variables (2BB-2F)
This approach uses the energy assets and the balance nodes as building blocks. In addition, it uses two positive flow variables to connect the nodes instead of the four variables used in the previous approach. The two flow variables are needed to model transmission or distribution network losses. However, gas flows with linepack cannot be modelled correctly with only two flow variables. Depending on the input data, the incoming and outgoing flow can have values simultaneously, especially when losses are not considered. In such cases, modellers typically use the net value as the transfer between the nodes. Alternatively, they may include a binary variable to prevent simultaneous incoming and outgoing flows, although this would convert the problem into a Mixed-Integer Programming problem.